OpenStudy (sims):

Look at triangle ABC: A right triangle ABC is shown with the right angle at B. The length of side AB is labeled as 14 mm, and the length of side BC is labeled as 48 mm. What is the length (in millimeters) of side AC of the triangle? 34 millimeters 50 millimeters 62 millimeters 70 millimeters

OpenStudy (sims):

Help please

OpenStudy (3mar):

If I may help?

OpenStudy (sims):

I got it wrong

OpenStudy (3mar):

 \[\large L=\sqrt{14^2+48^2}=\sqrt{2500}=50mm\] Done!

OpenStudy (sims):

I thought it was 70

OpenStudy (3mar):

No. You got it now. so the answer is ...#... ?

OpenStudy (sims):

Thank u

OpenStudy (3mar):

Honor is mine! Not at all! Thank you for learning!

OpenStudy (sims):

It Get it know

OpenStudy (3mar):

I am happy to hear that.

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (3mar):

I really have to go !!! sorry !!! Do you want anything for now?

OpenStudy (sims):

Your welcome

OpenStudy (3mar):

If you need help, I will help you as much as I know and as I can. Salam!